From GlueXWiki
Jump to: navigation, search

May 30, 2013 FDC meeting


  1. Production Construction Tracking
    • Spare package status
  2. Engineering (Bill)
    • Cooling system test
  3. Full electronics test (Beni)
  4. Other


Participants: Bill, Dave, Chris, Nick, Simon, Beni, Eugene, and Lubomir.


- Spare package second cell: It was tested over the weekend. There were once-per-night trips on Friday and Saturday, but from Sunday till Wednesday there were no trips even keeping for several hours higher HV (+2250/-500V). Interestingly, it turned out that sometimes when using the low-intensity Fe source (100-200 kHz rates on chamber) the dark current may increase significantly. We suspect this is again a glow discharge triggered by the source. However, with the time this effect was diminishing. Most of the sense wires of this old wire board (#3) have some spots that are not shiny at one side of the chamber covering an area of about 1/3 of the total, which may be causing the glow discharge. In any case this effect disappears if using say 2120/-400V instead of the nominal 2100/-500V, that has the same gas gain. Therefore, we conclude it is not worth replacing all the wires and this wire plane can be used.

- Spare package third and fourth cells: were installed on Wedneday with the help of Tina (plus Casey, Anatoly and Lubomir). Later it turned out there's a significant leak in the second cell at two places at the spacer ring. When adding the third and fourth cells we first need to take out the top cathode type-3 of the second cell and most likely the O-ring jumped out of the groove. The O-rings on this spacer were very old, used also for tests of wire frames #3 and #6 between Lexan sheets. The plan is first to test the two new cells and then to open the package and fix the O-ring. If there's nothing to be fixed on the new cells we will access the O-ring by lifting the top two-and-a-half cells together. Hopefully we can get our HV supply back on Friday to test the cells over the weekend. The plan is to open the package on Thursday after the collaboration meeting.


- Beni and Bill measured the max temperature on one of the card at the voltage regulator: no more than 43degC with the cooling system running. Beni: when turning the cooling system on, the temperature on the voltage regulator changes very fast indicating good thermal contact to the cooling tube.

- Bill once more on the SMC power: the chiller produces 2kW from the pump + 2kW from the compressor + max 2kW removed heat. Eugene: using 4kW to remove 2kW is something standard for such cooling systems.

- Beni started working on Linux interface to the SMC chiller using USB to RS485 adapter. His plan is to test the communication and make some basic control functions, that need to be discussed further.

Full electronics tests

- Beni can now read the new fADC125 modules; demonstrated it by reading three such modules. The basic firmware/software made by Cody and Bryan allows to read the raw data from the modules. The extraction of the amplitudes and timings is done in the ROCs. Even this will slow down the DAQ, it will be enough for the full package test with comics. Beni, however, has some issues with the TI synchronization with the modules. Bryan was helping him after the meeting.

- Right after the meeting we moved the third package from Blue Crab to EEL126. The package was connected to the gas and the pressure is OK. On Tuesday Casey will connect the package to the water chiller. The cabling will be done after the collaboration meeting.


- No meeting next week due to the collaboration meeting.